# references.bib

### v1

@@ 5.1 Algebraic convergence rate
To prove the result one needs to identify $\mu_{t}$ with $\mu_{t}^{0}$ and the supplementary regularity assumptions on $\mu_{t}^{\lambda}$ are sufficient for this purpose. Let us fix some $t\in[0,T]$ and note $u^{\lambda}_{t}=u^{\lambda}_{f}[\mu^{\lambda}_{t}]$ the minimizer in eq. 9, $\nu^{\lambda}_{t}\in\mathcal{M}(\Omega)$ the minimizer in eq. 18 s.t. $\frac{\mathrm{d}\nu^{\lambda}_{t}}{\mathrm{d}\mu^{\lambda}_{t}}=u^{\lambda}_{t}$ and $\alpha^{\lambda}_{t}\in L^{2}(\rho)$ the maximizer in eq. 19.
Then for every $\lambda>0$, since $\mu_{0}$ has bounded log-density we have by the flow-map representation in Proposition 4.1 that $\mu^{\lambda}_{t}$ has bounded log-density. Also, since $\bar{\nu}$ is positive with bounded log-density and $\Phi\star$ is injective, we have that $u^{\dagger}_{t}\coloneqq\frac{\mathrm{d}\bar{\nu}}{\mathrm{d}\mu^{\lambda}_{t}}$ is the unique solution to eq. 10. But then, by definition of the RKHS $\mathcal{H}$ we have that $\Phi^{\top}:L^{2}(\rho)\to\mathcal{H}$ is a partial isometry by Mercer’s theorem (Theorem A.1) and the assumption that $\partial f(\frac{\mathrm{d}\bar{\nu}}{\mathrm{d}\mu^{\lambda}_{t}})\in\mathcal{H}$ is equivalent to a source condition of the form in eq. 40. Hence by Lemma 5.1, the dual variable $\alpha^{\lambda}_{t}$ is bounded in $L^{2}(\rho)$, uniformly over $\lambda>0$, which implies that (up to extraction of a subsequence) $\Phi^{\top}\alpha^{\lambda}_{t}$ converges to some $h_{t}$ in $\mathcal{C}^{1}(\Omega)$.
Also for every $\lambda>0$, by the duality relations in eq. 20, we have $\frac{\mathrm{d}\nu^{\lambda}_{t}}{\mathrm{d}\mu^{\lambda}_{t}}=\partial f^{*}(\Phi^{\top}\alpha^{\lambda}_{t})$ and hence (recall $f(t)=|t|^{r}/(r-1)$ for some $r>1$) $\frac{\mathrm{d}\nu^{\lambda}_{t}}{\mathrm{d}\mu^{\lambda}_{t}}\to\partial f^{*}(h_{t})$ in $\mathcal{C}^{0}(\Omega)$. Since $\mathcal{L}^{\lambda}_{r}(\mu^{\lambda}_{t})\leq\mathcal{L}^{0}_{r}(\mu_{0})$ is bounded we have by eq. 18 that $\nu^{\lambda}_{t}\to\bar{\nu}$ narrowly and then for every $\varphi\in\mathcal{C}^{0}(\Omega)$:
@@
Since $f(t)=|t|^{r}/(r-1)$ for some $r>1$ the above equation is equivalent to eq. 38 which shows $\mu_{t}$ is the weak solution starting from $\mu_{0}$ of the ultra-fast diffusion equation eq. 36 according to Definition 4.2, that is $\mu_{t}=\mu^{0}_{t}$. ∎
The proof of the above Theorem 5.2 relies on the following result on solutions to inverse problems with nonlinear regularization [benning2018modern]. The following Lemma 5.1 is similar to [iglesias2018note, Prop. 3].
@@
Assume $f$ satisfies Assumption 2 and Assumption 1 holds. For $\mu\in\mathcal{P}(\Omega)$, let $u^{\dagger}\in L^{1}(\mu)$ be a solution of eq. 10. We say $u^{\dagger}$ satisfies a source condition if there exists $\alpha\in L^{2}(\rho)$ s.t.
@@
Then in this case, noting $\alpha^{\dagger}\in L^{2}(\mu)$ the $\alpha$ of minimal norm satisfying the above source condition, we have for every $\lambda>0$
@@
and integrating w.r.t. $\mu$ and using that $\int_{\Omega}(\Phi^{\top}\alpha^{\dagger})u^{\dagger}\mathrm{d}\mu=\langle\alpha^{\dagger},Y\rangle_{L^{2}(\rho)}$ we obtain:
$-\int_{\Omega}f^{*}(\Phi^{\top}\alpha^{\dagger})\mathrm{d}\mu+\langle\alpha^{\dagger},Y\rangle_{L^{2}(\rho)}\geq\int_{\Omega}f(u^{\dagger})\mathrm{d}\mu=\inf_{\Phi_{\mu}\cdot u=Y}\int_{\Omega}f(u)\mathrm{d}\mu.$
@@
$-\int_{\Omega}f^{*}(\Phi^{\top}\alpha^{\dagger})\mathrm{d}\mu+\langle\alpha^{\dagger},Y\rangle_{L^{2}(\rho)}\geq-\int_{\Omega}f^{*}(\Phi^{\top}\alpha)\mathrm{d}\mu+\langle\alpha,Y\rangle_{L^{2}(\rho)}.$
@@
$-\int_{\Omega}f^{*}(\Phi^{\top}\alpha^{\lambda})\mathrm{d}\mu+\langle\alpha^{\lambda},Y\rangle_{L^{2}(\rho)}-\frac{\lambda}{2}\|\alpha^{\lambda}\|^{2}_{L^{2}(\rho)}\geq-\int_{\Omega}f^{*}(\Phi^{\top}\alpha^{\dagger})\mathrm{d}\mu+\langle\alpha^{\dagger},Y\rangle_{L^{2}(\rho)}-\frac{\lambda}{2}\|\alpha^{\dagger}\|^{2}_{L^{2}(\rho)}.$
@@
$-\int_{\Omega}f^{*}(\Phi^{\top}\alpha^{\lambda})\mathrm{d}\mu+\langle\alpha^{\lambda},Y\rangle_{L^{2}(\rho)}-\frac{\lambda}{2}\|\alpha^{\lambda}\|^{2}_{L^{2}(\rho)}\geq-\int_{\Omega}f^{*}(\Phi^{\top}\alpha)\mathrm{d}\mu+\langle\alpha,Y\rangle_{L^{2}(\rho)}-\frac{\lambda}{2}\|\alpha\|^{2}_{L^{2}(\rho)}$
$\xrightarrow{\lambda\to 0+}$ $-\int_{\Omega}f^{*}(\Phi^{\top}\alpha^{0})\mathrm{d}\mu+\langle\alpha^{0},Y\rangle_{L^{2}(\rho)}\|^{2}_{L^{2}(\rho)}\geq-\int_{\Omega}f^{*}(\Phi^{\top}\alpha)\mathrm{d}\mu+\langle\alpha,Y\rangle_{L^{2}(\rho)},$
which shows $\alpha^{0}$ is also a maximizer of the dual problem eq. 19 when $\lambda=0$ and, as a consequence also satisfies the source condition eq. 40. But, by minimality of the norm of $\alpha^{\dagger}$ and by weak lower-semicontinuity of the norm we have:

### v2

@@ 5.1 Algebraic convergence rate
To prove the result one needs to identify $\mu_{t}$ with $\mu_{t}^{0}$ and the supplementary regularity assumptions on $\mu_{t}^{\lambda}$ are sufficient for this purpose. Let us fix some $t\in[0,T]$ and denote by $u^{\lambda}_{t}=u^{\lambda}_{f}[\mu^{\lambda}_{t}]$ the minimizer in Eq. 17, $\nu^{\lambda}_{t}\in\mathcal{M}(\Omega)$ the minimizer in Eq. 26 s.t. $\frac{\mathrm{d}\nu^{\lambda}_{t}}{\mathrm{d}\mu^{\lambda}_{t}}=u^{\lambda}_{t}$ and $\alpha^{\lambda}_{t}\in L^{2}(\rho)$ the maximizer in Eq. 27.
Then for every $\lambda>0$, since $\mu_{0}$ has bounded log-density we have by the flow-map representation in Proposition 4.1 that $\mu^{\lambda}_{t}$ has bounded log-density. Also, since $\bar{\nu}$ is positive with bounded log-density and $\Phi\star$ is injective, we have that $u^{\dagger}_{t}\coloneqq\frac{\mathrm{d}\bar{\nu}}{\mathrm{d}\mu^{\lambda}_{t}}$ is the unique solution to Eq. 18. But then, by the characterization of the RKHS $\mathcal{H}$ in Theorem 6, we have that $\Phi^{\top}:L^{2}(\rho)\to\mathcal{H}$ is a partial isometry and the assumption that $\partial f(\frac{\mathrm{d}\bar{\nu}}{\mathrm{d}\mu^{\lambda}_{t}})\in\mathcal{H}$ is equivalent to a source condition of the form Eq. 48. Hence by Lemma 5.1 the dual variable $\alpha^{\lambda}_{t}$ is bounded in $L^{2}(\rho)$, uniformly over $\lambda>0$, which implies that, up to extraction of a subsequence, $\Phi^{\top}\alpha^{\lambda}_{t}$ converges to some $h_{t}$ in $\mathcal{C}^{1}(\Omega)$.
Also for every $\lambda>0$, by the duality relations in Eq. 28, we have $\frac{\mathrm{d}\nu^{\lambda}_{t}}{\mathrm{d}\mu^{\lambda}_{t}}=\partial f^{*}(\Phi^{\top}\alpha^{\lambda}_{t})$ and hence — recalling that $f(t)=|t|^{r}/(r-1)$ for some $r>1$$\frac{\mathrm{d}\nu^{\lambda}_{t}}{\mathrm{d}\mu^{\lambda}_{t}}\to\partial f^{*}(h_{t})$ in $\mathcal{C}^{0}(\Omega)$. Since $\mathcal{L}^{\lambda}_{r}(\mu^{\lambda}_{t})\leq\mathcal{L}^{0}_{r}(\mu_{0})$ is bounded we have by Eq. 26 that $\nu^{\lambda}_{t}\to\bar{\nu}$ narrowly and then for every $\varphi\in\mathcal{C}^{0}(\Omega)$:
@@
Since $f(t)=|t|^{r}/(r-1)$ for some $r>1$ the above equation is equivalent to Eq. 46 which shows $\mu_{t}$ is the weak solution starting from $\mu_{0}$ of the ultra-fast diffusion equation Eq. 44 according to Definition 2, that is $\mu_{t}=\mu^{0}_{t}$. ∎
The proof of the above Theorem 5 relies on the following result on solutions to inverse problems with nonlinear regularization [benning2018modern]. The following Lemma 5.1 is similar to [iglesias2018note, Prop. 3].
@@
Let $f$ satisfy Assumption 2 and Assumption 1. For $\mu\in\mathcal{P}(\Omega)$, let $u^{\dagger}\in L^{1}(\mu)$ be a solution of Eq. 18. We say $u^{\dagger}$ satisfies a source condition if there exists $\alpha\in L^{2}(\rho)$ s.t.
@@
Then in this case, noting $\alpha^{\dagger}\in L^{2}(\rho)$ the $\alpha$ of minimal norm satisfying the above source condition, we have for every $\lambda>0$:
@@
and integrating w.r.t. $\mu$ and using that $\int_{\Omega}(\Phi^{\top}\alpha^{\dagger})u^{\dagger}\mathrm{d}\mu=\left<\alpha^{\dagger},Y\right>_{L^{2}(\rho)}$ we obtain:
$-\int_{\Omega}f^{*}(\Phi^{\top}\alpha^{\dagger})\mathrm{d}\mu+\left<\alpha^{\dagger},Y\right>_{L^{2}(\rho)}\geq\int_{\Omega}f(u^{\dagger})\mathrm{d}\mu=\inf_{\Phi_{\mu}\cdot u=Y}\int_{\Omega}f(u)\mathrm{d}\mu.$
@@
$-\int_{\Omega}f^{*}(\Phi^{\top}\alpha^{\dagger})\mathrm{d}\mu+\left<\alpha^{\dagger},Y\right>_{L^{2}(\rho)}\geq-\int_{\Omega}f^{*}(\Phi^{\top}\alpha)\mathrm{d}\mu+\left<\alpha,Y\right>_{L^{2}(\rho)}.$
@@
$-\int_{\Omega}f^{*}(\Phi^{\top}\alpha^{\lambda})\mathrm{d}\mu+\left<\alpha^{\lambda},Y\right>_{L^{2}(\rho)}-\frac{\lambda}{2}\|\alpha^{\lambda}\|^{2}_{L^{2}(\rho)}\geq-\int_{\Omega}f^{*}(\Phi^{\top}\alpha^{\dagger})\mathrm{d}\mu+\left<\alpha^{\dagger},Y\right>_{L^{2}(\rho)}-\frac{\lambda}{2}\|\alpha^{\dagger}\|^{2}_{L^{2}(\rho)}.$
@@
$-\int_{\Omega}f^{*}(\Phi^{\top}\alpha^{\lambda})\mathrm{d}\mu+\left<\alpha^{\lambda},Y\right>_{L^{2}(\rho)}-\frac{\lambda}{2}\|\alpha^{\lambda}\|^{2}_{L^{2}(\rho)}\geq-\int_{\Omega}f^{*}(\Phi^{\top}\alpha)\mathrm{d}\mu+\left<\alpha,Y\right>_{L^{2}(\rho)}-\frac{\lambda}{2}\|\alpha\|^{2}_{L^{2}(\rho)}$
$\xrightarrow{\lambda\to 0+}$ $-\int_{\Omega}f^{*}(\Phi^{\top}\alpha^{0})\mathrm{d}\mu+\left<\alpha^{0},Y\right>_{L^{2}(\rho)}\geq-\int_{\Omega}f^{*}(\Phi^{\top}\alpha)\mathrm{d}\mu+\left<\alpha,Y\right>_{L^{2}(\rho)},$
which shows $\alpha^{0}$ is also a maximizer of the dual problem Eq. 27 when $\lambda=0$ and, as a consequence, also satisfies the source condition Eq. 48. But, by minimality of the norm of $\alpha^{\dagger}$ and by weak lower semicontinuity of the norm we have: